find the area of an rectangle that has a perimeter of 20m and 10m
Answer by Earlsdon(6294)

Impossible!
If a rectangle has a perimeter of 20 meters, you cannot determine its area until you know its length and width. With a perimeter of 20 meters, its length and width could be: (...and these are just a few examples)
L = 9 and W = 1 P = 2(L+W) = 2(9+1) = 20, A = L*W = 9*1 = 9
L = 8 and W = 2 P = 2(L+W) = 2(8+2) = 20, A = L*W = 8*2 = 16
L = 6 and W = 4 P = 2(L+W) = 2(6+4) = 20, A = L*W = 6*4 = 24
Perhaps you meant to write:
"Find the area of a rectangle that has a length of 20m and a width of 10m"??
This can be done:
A = L*W
A = 20*10
A = 200 sq.meters.